X=842, n=1200, interval estimate 95%

1
Expert's answer
2022-05-02T15:54:06-0400

The sample proportion is computed as follows, based on the sample size "N = 1200" and the number of favorable cases "X = 842:"

"\hat{p}=\dfrac{X}{N}=\dfrac{842}{1200}\approx0.701667"

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(0.701667-1.96\sqrt{\dfrac{0.701667(1-0.701667)}{1200}},"

"0.701667+1.96\sqrt{\dfrac{0.701667(1-0.701667)}{1200}})"

"=(0.615,0.788)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.615 < p < 0.788," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.615, 0.788)."